# Two liquid states of distinguishable helium-4: the existence of another non-superfluid frozen by heating

### v2

@@ III.4 Quantum wavelength and expansion factor
To quantify the spatial extension of atomic necklaces, i.e., the degree of quantum dispersion, we evaluated quantum wavelength [45, 21] $\lambda_{\rm{quantum}}$ defined as the square root of the imaginary-time mean square displacements (IMSD) [53] at its maximum. Namely, the IMSD at imaginary time $\tau_{n}=n\beta\hbar/N_{\rm{b}}$ is defined as
@@
while the quantum wavelength $\lambda_{\rm{quantum}}$ is
@@
Surprisingly, these values are nearly equal to those of the two detailed states of supercritical fluid (SCF) 4He ($-0.29\leq\chi\leq-0.26$ for liquid or liquid-like state, $-0.49\leq\chi\leq-0.46$ for the gas or gas-like state) [45]. In this sense, the LQD-HQD boundary in the $\lambda$-$T$ plot is similar to the Widom line [45] that separates the liquid-like and gas-like SCF states by the $C_{P}$ maxima. However, as seen in Sec. III.2, there are no traces of such $C_{P}$ maxima in the present case.
@@
Figure 8: The expansion factor vs (a) temperature and (b) quantumness parameter. The colors of the symbols are the same as those displayed in Fig. 1: LQDL (blue), LQDA (cyan), HQDL (red), and HQDA (magenta). In (b), when there is a relation between the de Brogle thermal wavelength and the closest distance, $\lambda_{\rm{dB}}\geq\sigma$, we get $\Lambda^{*}\geq 1/\sqrt{2\pi}=0.40$. The yellow background shown in (b) denotes a possible universal curve for the LQD states.
The state transition can be detected more clearly by use of expansion factor $\alpha:=\lambda_{\rm{quantum}}/\lambda_{\rm{quantum}}^{(0)}$ where the denominator is the quantum wavelength of a three-dimensional free particle, $\lambda_{\rm{quantum}}^{(0)}=(\hbar/2)\sqrt{3\beta/m}=\sqrt{3/8\pi}\lambda_{\rm{dB}}$ [21]. We can regard $\alpha$ as being equivalent to expansion factor in polymer physics, i.e., $\alpha\approx R_{\rm{g}}/{{R_{\rm{g}}}^{\rm{(free)}}}$, where $R_{\rm{g}}^{\rm{(free)}}$ is the radius of gyration of classical free ring polymer. Figure 8(a) shows the temperature dependence of $\alpha$. In every isobar there exists a minimum or the tipping point locating at 0.2-0.6 K. As the temperature lowers, $\alpha$ decreases in the LQDL and LQDA, whereas it reversely increases in the HQDL and HQDA. Namely, each minimum point in this $\alpha$-$T$ plot is the transition point between a LQD state and an HQD state. A positive (negative) slope of the $\alpha$-$T$ curve corresponds to a LQD state (an HQD state), respectively.
Figure 8(b) shows the replot of $\alpha$ shown in Fig. 8(a) as a function of quantumness parameter $\Lambda^{*}$. Here the interatomic closest distance $\sigma$ was taken as the distance of the first peak of $g_{\rm{bb}}$. In this figure, $\Lambda^{*}$ of the HQD states are higher than those at tipping points. Interestingly, the $\Lambda^{*}$ at the tipping points ($1\lesssim{\Lambda^{*}}\lesssim 3$) is in good agreement with $\Lambda^{*}$ in the the quantum binary Lennard-Jones system ($1\lesssim\Lambda^{*}\lesssim 2$[8, 9]. Seemingly, in Fig. 8(b), plotted points of LQD states tend to gather on a single curve which is possibly located at yellow background region. This suggests the universality of the relation between $\alpha$ and $\Lambda^{*}$ of the LQD states, whereas the data points of the HQD states are remarkably deviated from it.
@@ III.5 Apparent self-diffusion coefficient
Finally, we provide the $D$-$\Lambda^{*}$ relation in Fig. SA-12 in the Supplementary Material. Compared with Fig.  9, the plotted points of LQD states seemingly tend to be located on a single line. This suggests the possible universality of the is $D$-$\Lambda^{*}$ relation underlying the LQD states.
@@ III.6 Identification of states
On the basis of the results presented in Secs. III.2III.5, we identified the state at each $P$-$T$ point by following the criteria:
@@
The conclusion of such identification of the system state was summarized in Fig. 1.
@@ III.7 Quantum tunneling of atoms
Among the graphs at 50 bar, the HQDA at 0.1 K and 50 bar has the largest population of negative $K_{i}$. This seems to be in harmony with the the double-well tunneling model to explain the linear temperature-dependence of heat capacity, $C_{V}{\sim}{T}$, of quantum glasses near $T\sim 0$ [10, 11]. The existence of a low concentration of tunneling atoms in HQDL at lowermost temperatures likely suggests a possible deviation from the heat capacity in the phonon regime, $C_{V}{\sim}T^{3}$, in this liquid state near $T{\sim}0$ if it exists.
@@ IV ANALYSIS BASED ON THE STATE DIAGRAM
In Fig. 1, the presumed melting curve drawn in parallel with the glass transition line is downward convex and have a minimum at $T_{\rm{min}}\cong 0.4$ K and $P_{\rm{min}}\cong 25$ bar. In general, a melting line with an extremum can be related to the occurrence of liquid polyamorphism [35]; two liquid states may exist at higher and lower temperatures than $T_{\rm{min}}$. In accordance with this, the LQDL-HQDL boundary line emanates from the vicinity of the minimum of the presumed melting curve. This boundary line is located in parallel with the experimental $\lambda$ transition line, and is shifted low by about 1.8 K. Boninsegni et al. presumed the phase diagram of distinguishable 4He, where the solid-liquid boundary was also downward convex [6]. In fact, the experimental melting curve of 4He (3He) has a shallow minimum at $T_{\rm{min}}=0.8$ K and $P_{\rm{min}}=26$ bar for 4He (0.3 K and 29 bar for 3He). The melting line in Fig. 1 is much more emphasized convex than the real system.
When a state point on the $P$-$T$ plane moves across the HQDL-LQDA boundary line within the range of $0.2\leq{T}\leq 0.3$ K and $40\leq{P}\leq 50$ bar, melting by cooling (freezing by heating) or exothermic melting (endothermic freezing) occurs; these transitions are alternatively called inverse melting (freezing) [36, 37, 38, 41, 40, 58, 39]. Thus, the state diagram evidences the occurrence of isobaric inverse melting (freezing). In order to confirm it, however, such inverse freezing of HQDL by heating was examined in this study, and the results are shown in Appendix; we succeeded in observing the inverse freezing in this test simulation.
@@
Under the assumption that Eq. (13) approximately holds for the glass transition line in a crude approximation, we can further proceed analysis of the results. Namely, we reread Eq. (13) as a pseudo-Clapeyron equation, approximately denoting the slope of glass transition line,
@@
where $P_{\rm{g}}$ and $T_{\rm{g}}$ are the glass transition pressure and temperature, respectively, while ${\Delta}V_{1\rightarrow{2}}$ and ${\Delta}H_{1\rightarrow{2}}$ are the volume and enthalpy change over a finite temperature width $\Delta{T}_{1\rightarrow{2}}$ riding on the glass transition line, $T_{1}<T_{\rm{g}}<T_{2}$, respectively. We presume that this crude relation is never far from reality, because the difference between thermodynamic property $X$ of a glass and the crystal is, in general, small enough and the slopes of the glass transition line and of the melting curve are nearly equal. A similar treatment was carried out in our previous study where the pseudo-Clapeyron equation was applied to the Widom line, which we regarded as the pseudo-boiling line separating the gas-like and liquid-like SCF [45]. In Fig. 3, we can see ${\Delta}V=V_{\rm{LQDA,0.3K}}-V_{\rm{HQDL,0.2K}}<0$ at $0.2\leq{T_{g}}\leq 0.3$ K for the isobars of 40 and 50 bars, while in Fig. 4 enthalpy change with increasing $T$ is always positive, ${\Delta}H>0$, for any isobars. Accordingly, Eq. (14) yields ${dP_{\rm{m}}}/{dT}\cong{dP_{\rm{g}}}/{dT}<0$, coinciding with negative slope of the left half ($T<T_{\rm{m}}$) of the glass transition line and presumed melting line in Fig. 1. Therefore, the presumed equilibrium crystal phase overlapping on the convex region of LQDA at 0.3 K and 40-50 bars should have larger entropy than the HQDL at 0.2 K and 40-50 bars. On the other hand, the slope of the right half ($T>T_{\rm{m}}$) of the glass transition line and the presumed melting line are positive, $dP_{\rm{m}}/dT\cong{dP_{\rm{g}}}/{dT}>0$, indicating the normal melting (melting by heating). This is because we can see, in the range of $3.30.4{\leq}T{\leq}$ 3.3 K and $50040{\leq}P{\leq}$ 500 bar, ${\Delta}H_{{\rm{solid(LQDA)}}\rightarrow{\rm{liquid(LQDL)}}}>0$ (Fig. 4) and ${\Delta}V_{{\rm{solid(LQDA)}}\rightarrow{\rm{liquid(LQDL)}}}>0$ (Fig. 3). Thus, the analysis based on the pseudo-Clapeyron equation evidences the downward convex curves of the glass transition line and the presumed melting line shown in Fig. 1.
@@
Figure 12: The state diagram of distinguishable 4He obtained from isothermal compression over wider $P$-$T$ range. LQDL (blue): low quantum dispersion liquid; HQDL (red): high quantum dispersion liquid; LQDA (cyan): low quantum dispersion amorphous solid (metastable); HQDA (magenta): high quantum dispersion amorphous solid (metastable). “LQD fluid‘ includes LQDL, gas (in LQD state), and supercritical fluid (in LQD state) existing beyond the critical point (red star). Black solid lines are the experimental phase boundary lines of real 4He. Experimentally, He-I and He-II (gray slant stripe) exist at higher and lower temperatures than $\rm{\lambda}$ transition line (black), respectively, while the crystalline solid phase exists at pressures higher than the melting curve (black). The open symbols are the points explored by the isothermal compression simulation in our previous study [21]. The Kim-Chan boundary line is cited from Ref. 32.
@@ V DISCUSSIONS
V.1 Relevance of HQDL to real non-superfluid states
Figure 13: The state diagram of distinguishable 4He compared with the states of confined real 4He systems. LQDL (blue): low quantum dispersion liquid; HQDL (red): high quantum dispersion liquid; LQDA (cyan): low quantum dispersion amorphous solid (metastable); HQDA (magenta): high quantum dispersion amorphous solid (metastable). The Kim-Chan phase boundary is cited from Ref. 32. The phase boundary lines of real 4He doped in Gelsil (25 Å) [62], folded sheets meso-porous materials (FSM) (28 Å), [63] and MCM-41 (core diameter 47 Å) [64], are drawn together. “$T_{\rm{C}}$, FSM”: superfluid onset temperature of the FSM system; “$T_{\rm{B}}$, FSM”: rapid growth temperature of the superfluid fraction of the FSM system [63]; “$T_{\rm{C}}$, Gelsil”: superfluid transition temperature of the Gelsil system; “$T_{\rm{B}}$, Gelsil”: localized BEC (LBEC) temperature of the Gelsil system [62]. Experimentally, the BEC-like low-entropy state exists in gray slant stripe region [63], while the LBEC exists in gray polka-dot region [62].
Recent experiments targeting confined 4He revealed that the non-superfluid states, unlike He-I and He-II, exist at temperatures lower than the ${\rm{\lambda}}$-transition line of the bulk system [65, 66, 67, 62, 68, 69, 70, 63, 71, 64, 72]. Under such confined conditions, the off-diagonal long-range Bosonic correlation is hindered by the the wall of glass substrate pores, so that Bosonic permutation effect should partially be suppressed [68, 62]. Figure  13 shows the comparison of our state diagram with that of confined 4He in three kinds of substrates. Shirahama and co-workers revealed that the localized BEC (LBEC) state of 4He confined in Gelsil exists in the region between the boundaries of “$T_{\rm{B}}$, Gelsil” and “$T_{\rm{C}}$, Gelsil” [65, 66, 67, 62, 68]. For 4He confined in the folded sheets meso-porous materials (FSM) (28 Å), Taniguchi et al. reported that the state at temperatures between “$T_{\rm{B}}$, FSM” and “$T_{\rm{C}}$, FSM” was not He-I nor He-II but a BEC-like state with low entropy [63]. We can see that the experimental state boundaries have the same negative slopes as the HQDL-LQDA boundary line. Notably, some part of the $P$-$T$ region of HQDL overlaps with those of the LBEC and the low-entropy BEC-like states. One sees that the LBEC region overlaps with the upward convex region of HQDL (in the range of $18\lesssim{P}\lesssim{35}$ bar and $T\lesssim{0.35}$ K) causing inverse freezing. Since the Bosonic long-range correlation is considerably suppressed in the confined 4He systems, we speculate that LBEC and BEC-like non-superfluid states have some relevance to the HQDL which obeys Boltzmann statistics.
Surprisingly, in Fig. 1, the HQDL-LQDA boundary emanating from the point at $T=0.3$ K and $P=26$ bar traces well the Kim-Chan boundary [32]. This boundary was interpreted as an evidence of supersolid which might exist at lower-temperature side of the line, on the basis of the detection of non-classical moment of inertia by the torsional oscillator experiment [32]. However, nowadays the existence of defect-less crystalline supersolid phase is almost denied, while this anomaly is attributed to some disorder due to dislocation, grain boundaries, glassy domains, and 3He impurities [73]. In such a narrow domain unlike a bulk, the long-range atomic permutation should partially be suppressed so that the BEC should not completely be fulfilled. It is noted again that the HQDL has the upward convex $P$-$T$ region at $T\lesssim 0.3$ K and $P\gtrsim 25$ bar, which coincides with reported supersolid region. Therefore, we speculate that some gas-like property can be observed experimentally at the supersolid $P$-$T$ region which is, in the Boltzmann regime, the upward convex region of HQDL. This would be possible when an observation is targeted on non-bulk domains (i.e., confined systems and disordered domains included in the bulk solid) which obeys suppressed Bose statistics (or semi-Boltzmann statistics).
V.2 On state transition

### v3

@@ III.3 Radial distribution functions
We show the $g_{\rm{cc}}$ at 1 bar with $N_{\rm{b}}=500$, 1500, and 2000 in Fig. SA-4 in the Supplementary Material. For each of $N_{\rm{b}}$, we can see that the LQD-HQD structural transition occurred at 0.5-0.7 K; the structure of HQD state is characterized by non-zero distribution of $g_{\rm{cc}}$ at the distances shorter than the first peak. Consequently, the occurrence of the LQD-HQD transition is not artificially caused by the insufficiency of $N_{\rm{b}}$ but attributed to an essential property of the system.
@@ III.4 Quantum wavelength and expansion factor
To quantify the spatial extension of atomic necklaces, i.e., the degree of quantum dispersion, we evaluated quantum wavelength [46, 21] $\lambda_{\rm{quantum}}$ from the square root of the imaginary-time mean square displacements (IMSD) [54]. Namely, the IMSD at imaginary time $\tau_{n}=n\beta\hbar/N_{\rm{b}}$ is
@@
from which we estimated the quantum wavelength $\lambda_{\rm{quantum}}$ defined as
@@
Surprisingly, these values are nearly equal to those of the two detailed states of supercritical fluid (SCF) 4He ($-0.29\leq\chi\leq-0.26$ for liquid or liquid-like state, $-0.49\leq\chi\leq-0.46$ for the gas or gas-like state).[46] In this sense, the LQD-HQD boundary in the $\lambda$-$T$ plot is similar to the Widom line [46] that separates the liquid-like and gas-like SCF states by the $C_{P}$ maxima. However, as seen in Sec. III.2, there are no traces of such $C_{P}$ maxima in the present case.
@@
Figure 8: The expansion factor vs (a) temperature and (b) quantumness parameter. The colors of the symbols are the same as those displayed in Fig. 1: LQDL (blue), LQDA (cyan), HQDL (red), and HQDA (magenta). In (b), when there is a relation between the de Broglie thermal wavelength and the closest distance, $\lambda_{\rm{dB}}\geq\sigma$, we get $\Lambda^{*}\geq 1/\sqrt{2\pi}=0.40$. The yellow background shown in (b) denotes a possible universal curve for the LQD states.
The state transition can be detected more clearly by use of expansion factor $\alpha:=\lambda_{\rm{quantum}}/\lambda_{\rm{quantum}}^{(0)}$ where the denominator is the quantum wavelength of a three-dimensional free particle, $\lambda_{\rm{quantum}}^{(0)}=(\hbar/2)\sqrt{3\beta/m}=\sqrt{3/8\pi}\lambda_{\rm{dB}}$ [21]. We can regard $\alpha$ as being equivalent to expansion factor in polymer physics, i.e., $\alpha\approx R_{\rm{g}}/{{R_{\rm{g}}}^{\rm{(free)}}}$, where $R_{\rm{g}}^{\rm{(free)}}$ is the radius of gyration of classical free ring polymer. Figure  8(a) shows the temperature dependence of $\alpha$. In every isobar there exists a minimum or the tipping point locating at 0.2-0.6 K. As the temperature lowers, $\alpha$ decreases in the LQDL and LQDA, whereas it reversely increases in the HQDL and HQDA. Namely, each minimum point in this $\alpha$-$T$ plot is the transition point between an LQD state and an HQD state. A positive (negative) slope of the $\alpha$-$T$ curve corresponds to an LQD state (an HQD state), respectively.
Figure 8(b) shows the replot of $\alpha$ shown in Fig. 8(a) as a function of quantumness parameter $\Lambda^{*}$. Here the interatomic closest distance $\sigma$ was taken as the distance of the first peak of $g_{\rm{bb}}$. In this figure, $\Lambda^{*}$ of the HQD states are higher than those at tipping points. Interestingly, the $\Lambda^{*}$ at the tipping points ($1\lesssim{\Lambda^{*}}\lesssim 3$) is in good agreement with $\Lambda^{*}$ in the quantum binary Lennard-Jones system ($1\lesssim\Lambda^{*}\lesssim 2$) [8, 9]. Seemingly, in Fig. 8(b), plotted points of LQD states tend to gather on a single curve which is possibly located at yellow background region. This suggests the universality of the relation between $\alpha$ and $\Lambda^{*}$ of the LQD states, whereas the data points of the HQD states are remarkably deviated from it.
We comment on the bead number dependence of $\lambda_{\rm{quantum}}$ and $\alpha$. The examination of this dependence was examined by short CMD runs at 1 bar, and the graphs are shown in Fig. SA-3 in the Supplementary Material. In Fig. SA-3(a), clearly, remarkable increase of $\lambda_{\rm{quantum}}$ with lowering temperature toward 0.1 K is observed for every $N_{\rm{b}}$. Further, in Fig. SA-3(b) the reentrant temperature-dependence of $\alpha$ passing the tipping points is recognized for all $N_{\rm{b}}$ conditions. Therefore, the occurrence of the LQD-HQD transition is essential and is not an artifact.
@@ III.5 Apparent self-diffusion coefficient
Finally, we provide the $D$-$\Lambda^{*}$ relation in Fig. SA-14 in the Supplementary Material. Compared with Fig.  9, the plotted points of LQD states seemingly tend to be located on a single line. This suggests the possible universality of the $D$-$\Lambda^{*}$ relation underlying the LQD states.
@@ III.6 Identification of states
On the basis of the results presented in Secs. III.2III.5, we identified the state at each $P$-$T$ point by following the criteria (1)-(3):
@@
The conclusion of such identification of the system state was summarized in Fig. 1. Here, after plotting the identified points on the $P$-$T$ plane, the boundary lines were drawn passing between two adjacent points representing the two different states. In this figure, we discarded the data points at $T<0.1$ K from the plot because we cannot deny possible artifact due to the insufficiency of the number of beads ($N_{\rm{b}}=500$) at these lowermost temperatures.
@@ III.7 Quantum tunneling of atoms
Among the graphs at 50 bar, the HQDA at 0.1 K and 50 bar has the largest population of negative $K_{i}$. This seems to be in harmony with the double-well tunneling model to explain the linear temperature-dependence of heat capacity, $C_{V}{\sim}{T}$, of quantum glasses near $T\sim 0$ [10, 11]. The existence of a low concentration of tunneling atoms in HQDL at lowermost temperatures likely suggests a possible deviation from the heat capacity in the phonon regime, $C_{V}{\sim}T^{3}$, in this liquid state near $T{\sim}0$ if it exists.
@@ IV ANALYSIS BASED ON THE STATE DIAGRAM
In Fig. 1, the presumed melting curve drawn in parallel with the glass transition line is downward convex and has a minimum at $T_{\rm{min}}\cong 0.4$ K and $P_{\rm{min}}\cong 25$ bar. In general, a melting line with an extremum can be related to the occurrence of liquid polyamorphism [35]; two liquid states may exist at higher and lower temperatures than $T_{\rm{min}}$. In accordance with this, the LQDL-HQDL boundary line emanates from the vicinity of the minimum of the presumed melting curve. This boundary line is located in parallel with the experimental $\lambda$ transition line, and is shifted low by about 1.8 K. Boninsegni et al. presumed the phase diagram of distinguishable 4He, where the solid-liquid boundary was also downward convex [5]. In fact, the experimental melting curve of 4He (3He) has a shallow minimum at $T_{\rm{min}}=0.8$ K and $P_{\rm{min}}=26$ bar for 4He (0.3 K and 29 bar for 3He). The melting line in Fig. 1 is much more emphasized convex than the real system.
When a state point on the $P$-$T$ plane moves across the HQDL-LQDA boundary line within the range of $0.2\leq{T}\leq 0.3$ K and $40\leq{P}\leq 50$ bar, melting by cooling (freezing by heating) or exothermic melting (endothermic freezing) occurs; these transitions are alternatively called inverse melting (freezing) [36, 37, 38, 41, 40, 59, 39]. Thus, the state diagram evidences the occurrence of isobaric inverse melting (freezing). In order to confirm it, however, such inverse freezing of HQDL by heating was examined in this study, and the results are shown in Appendix A; we succeeded in observing the inverse freezing in this test simulation.
@@
We further proceed analysis of the results by rereading Eq. (13) as a pseudo-Clapeyron equation which approximately holds for the slope of glass transition line [62],
@@
where $P_{\rm{g}}$ and $T_{\rm{g}}$ are the glass transition pressure and temperature, respectively, while ${\Delta}V_{1\rightarrow{2}}$ and ${\Delta}H_{1\rightarrow{2}}$ are the volume and enthalpy change over a finite temperature width $\Delta{T}_{1\rightarrow{2}}$ riding on the glass transition line, $T_{1}<T_{\rm{g}}<T_{2}$, respectively. As shown in Ref. 62, this approximated relation is never far from reality, because the difference between thermodynamic property $X$ of a glass and the crystal is, in general, small enough and the slopes of the glass transition line and of the melting curve are nearly equal. A similar treatment was carried out in our previous study where the pseudo-Clapeyron equation was applied to the Widom line, which we regarded as the pseudo-boiling line separating the gas-like and liquid-like SCF [46]. In Fig. 3, we can see ${\Delta}V=V_{\rm{LQDA,0.3K}}-V_{\rm{HQDL,0.2K}}<0$ at $0.2\leq{T_{g}}\leq 0.3$ K for the isobars of 40 and 50 bars, while in Fig. 4 enthalpy change with increasing $T$ is always positive, ${\Delta}H>0$, for any isobars. Accordingly, Eq. (14) yields ${dP_{\rm{m}}}/{dT}\cong{dP_{\rm{g}}}/{dT}<0$, coinciding with negative slope of the left half ($T<T_{\rm{min}}$) of the glass transition line and presumed melting line in Fig. 1. Therefore, the presumed equilibrium crystal phase overlapping on the convex region of LQDA at 0.3 K and 40-50 bars should have larger entropy than the HQDL at 0.2 K and 40-50 bars. On the other hand, the slope of the right half ($T>T_{\rm{min}}$) of the glass transition line and the presumed melting line are positive, $dP_{\rm{m}}/dT\cong{dP_{\rm{g}}}/{dT}>0$, indicating the normal melting (melting by heating). This is because we can see, in the range of $3.30.4{\leq}T{\leq}$ 3.3 K and $50040{\leq}P{\leq}$ 500 bar, ${\Delta}H_{{\rm{solid(LQDA)}}\rightarrow{\rm{liquid(LQDL)}}}>0$ (Fig. 4) and ${\Delta}V_{{\rm{solid(LQDA)}}\rightarrow{\rm{liquid(LQDL)}}}>0$ (Fig. 3). Thus, the analysis based on the pseudo-Clapeyron equation evidences the downward convex curves of the glass transition line and the presumed melting line shown in Fig. 1.
@@
Figure 12: The state diagram of distinguishable 4He obtained from isothermal compression over wider $P$-$T$ range. LQDL (blue): low quantum dispersion liquid; HQDL (red): high quantum dispersion liquid; LQDA (cyan): low quantum dispersion amorphous solid (metastable); HQDA (magenta): high quantum dispersion amorphous solid (metastable). “LQD fluid” includes LQDL, gas (in LQD state), and supercritical fluid (in LQD state) existing beyond the critical point (red star). Black solid lines are the experimental phase boundary lines of real 4He. Experimentally, He-I and He-II (gray slant stripe) exist at higher and lower temperatures than $\rm{\lambda}$ transition line (black), respectively, while the crystalline solid phase exists at pressures higher than the melting curve (black). The open symbols are the points explored by the isothermal compression simulation in our previous study [21]. The Kim-Chan boundary line is cited from Ref. 32.
As described in Sec. I, the present system should be liquid at $T=0$. Since this should not violate the third law of thermodynamics, the system must be in a single state without degeneration [6, 63]. At $T=0$, therefore, the liquid phase must have zero entropy equal to the crystal phase existing at higher pressures, so that the presumed melting line should be horizontal owing to the Clausius-Clapeyron equation, Eq. (13). Although this line was drawn down to 0.1 K in Fig. 1, at present, it is unknown whether this becomes horizontal even at non-zero temperatures such as 0.01 or 0.001 K before reaching $T=0$. This issue would be revealed in a future work, together with exploring the possibility of another liquid polymorph at such lower temperatures.
@@ V DISCUSSIONS
V.1 On state transition